### v1

# On the Geometric Coherence of Global Aggregation in Federated GNN

Abstract
Federated Learning (FL) enables distributed training across multiple clients without centralized data sharing, while Graph Neural Networks (GNNs) model relational data through message passing. In federated GNN settings, client graphs often exhibit heterogeneous structural and propagation characteristics. When standard aggregation mechanisms are applied to such heterogeneous updates, the global model may converge numerically while exhibiting degraded relational behavior.Our work identifies a geometric failure mode of global aggregation in Cross- Domain Federated GNNs. Although GNN parameters are numerically represented as vectors, they encode relational transformations that govern the direction, strength, and sensitivity of information flow across graph neighborhoods. Aggregating updates originating from incompatible propagation regimes can therefore introduce destructive interference in this transformation space.This leads to loss of coherence in global message passing. Importantly, this degradation is not necessarily reflected in conventional metrics such as loss or accuracy.To address this issue, we propose GGRS (Global Geometric Reference Structure), a server-side framework that regulates client updates prior to aggregation based on geometric admissibility criteria. GGRS preserves directional consistency of relational transformations as well as maintains diversity of admissible propagation subspaces. It also stabilizes sensitivity to neighborhood interactions, without accessing client data or graph topology. Experiments on heterogeneous GNN-native, Amazon Co-purchase datasets demonstrate that GGRS preserves global message-passing coherence across training rounds by highlighting the necessity of geometry-aware regulation in federated graph learning.
keywords:
[…]
Article Info
 {history}
Received Feb 04, 2026
Revised month dd, yyyy
Accepted month dd, yyyy
Corresponding Author:
Chethana Prasad Kabgere
Department of Computer Science, Research Scholar, PES University, Karnataka, India
Email: chethana1999@gmail.com
    ntroduction
Federated Learning (FL) addresses distributed optimization problems in which a shared model is trained across multiple decentralized clients under communication and data-access constraints [1]. At each communication round, locally optimized parameter updates are aggregated at a global server to form a new global model, which is subsequently broadcast to clients for further local optimization.Whearas, whearas Cross-domain federated machine learning [4] refers to a federated setting in which participating clients hold data originating from distinct domains or distributions, rather than being non-IID partitions of a single dataset. In this regime, each client’s local data distribution may differ substantially in feature statistics, label semantics, or structural characteristics, leading to stronger heterogeneity than conventional FL. The goal remains to learn a shared global model via decentralized optimization, but aggregation becomes more challenging due to domain-induced misalignment across client updates. Classical aggregation schemes such as Federated Averaging (FedAvg) implicitly assume that client updates lie in a common parameter space where linear combination preserves model semantics.GNN[5] fundamentally violate this assumption. A GNN does not merely define a parameterized predictor, but rather a parameterized message-passing operator acting on graph-structured domains [5, 6]. For a fixed architecture, learnable parameters induce a family of linear and nonlinear operators that govern how node representations are updated as a function of neighborhood structure, aggregation depth, and interaction strength. Consequently, parameter updates in GNNs correspond to perturbations of an operator that determines the geometry of information flow over the graph, rather than independent numerical adjustments.
In federated GNN[2] settings, each client optimizes this operator with respect to its own graph distribution. Let $\mathcal{G}_{k}\sim\mathcal{D}_{k}$ denote the graph distribution observed by client $k$. Differences in topology, degree distribution, sparsity, and homophily across $\{\mathcal{D}_{k}\}$ induce structurally distinct optimal message-passing operators, even under a shared task objective [7]. Local training therefore shapes parameters toward client-specific propagation regimes, such as strong local aggregation, smooth long-range diffusion, or sparse selective propagation. These regimes correspond to distinct regions in the space of admissible operators.Standard global aggregation ignores this structure.We observe that by linearly averaging parameter updates, the server constructs a global parameter vector that is not guaranteed to correspond to a coherent message-passing operator for any of the underlying graph distributions. From an operator-theoretic perspective, this procedure amounts to mixing perturbations of distinct operators without enforcing compatibility of their dominant directions, invariant subspaces, or sensitivity profiles. Hence,the resulting global model may therefore exhibit operator degeneration, in which the induced message-passing transformation loses expressive capacity despite numerical convergence of parameters.This degeneration directly affects predictions. Inference in GNNs relies on the repeated application of the learned message-passing operator to propagate information across graph neighborhoods. We see that when aggregation disrupts the operator’s geometric structure, the global model can exhibit diminished sensitivity to graph perturbations, a collapse of effective propagation depth, or excessive contraction of node representations. These effects are closely related to over-smoothing and over-squashing phenomena studied in non-federated GNNs [8, 9]. Importantly, such failures may not immediately manifest in task loss or accuracy, as the model can still fit marginal label distributions while relying increasingly on non-relational cues.The impact of geometric degradation is amplified by the iterative nature of FL.
So in the further process, the global model produced by the server defines the initialization for all clients in the subsequent round. If this model encodes a distorted message-passing operator, local optimization must first compensate for this distortion before adapting to the local graph structure. This induces corrective gradients that are misaligned across clients, increasing update variance and amplifying destructive interference at the next aggregation step. Over multiple rounds, this feedback loop leads to progressive destabilization of the global operator, even as standard convergence criteria are satisfied.Existing approaches to federated GNN learning primarily address heterogeneity through personalization, client clustering, or adaptive weighting strategies[10, 2]. While effective in mitigating performance degradation, these methods do not explicitly characterize aggregation as an operation on message-passing operators, nor do they impose constraints that preserve the geometric integrity of the induced transformations. As a result, failure modes arising from operator-level incompatibility rather than optimization error remain unaddressed.
Our work adopts a geometric perspective on federated GNN aggregation, viewing client updates as perturbations in the space of message-passing operators. From this viewpoint, effective aggregation requires preserving structural properties such as directional alignment of operator updates, consistency of dominant subspaces, and stability of sensitivity to neighborhood structure. This formulation motivates the development of server-side mechanisms that regulate aggregation to maintain operator coherence across heterogeneous clients, thereby preventing silent degradation of global message-passing capacity in federated GNNs.
Figure 1: Geometric observations in federated GNN.
1.1  Geometric Observations in Federated GNN
From our study, we have summarized a set of fundamental observations regarding the geometric behavior of GNN under federated aggregation. These observations are not specific to a particular architecture or optimization strategy, but arise from the interaction between graph-dependent message passing and parameter aggregation across heterogeneous clients.Our first observation is Parameters as Structured Transformations: Although model parameters are numerically represented as vectors or matrices, learning behavior does not arise from individual scalar values but from coordinated changes across parameter groups. In GNNs, parameters jointly define structured transformations that govern how information propagates across graph neighborhoods. Consequently, the semantic meaning of a parameter update is determined by its direction and interaction with other parameters, rather than by its magnitude alone. In our next inference, we saw that GNNs Learn Message-Passing Operators: A GNN induces a parameterized message-passing operator whose repeated application updates node representations as a function of graph connectivity. In a fixed architecture, parameters specify how neighborhood information is aggregated, how far information propagates, and how sensitive node embeddings are to structural perturbations. Parameter updates therefore correspond to perturbations of this operator rather than independent numerical adjustments.Next we see,Graph-Dependent Operator Geometry: We know that the operator induced by a GNN is graph-dependent. The same parameter configuration can induce distinct transformations when applied to graphs with different adjacency structures or feature distributions. As a result, the Jacobian of the GNN with respect to its parameters depends on the underlying graph, implying that parameter updates computed on different clients generally inhabit distinct operator tangent spaces.Across the literature survey, we observed some implicit Assumptions of Federated Averaging. In Standard federated aggregation mechanisms implicitly assume that client updates are commensurate in a shared Euclidean parameter space, such that linear combination preserves model semantics. This assumption is violated in federated GNNs, where client updates encode perturbations of graph-dependent operators. Linear aggregation therefore mixes updates without enforcing compatibility of their induced transformations. Finally, we assessed the Numerical Convergence Versus Geometric Coherence. Federated training can exhibit numerical convergence in terms of loss reduction or gradient norms while simultaneously degrading the geometric coherence of the message-passing operator. Such degradation may manifest as diminished sensitivity to neighborhood structure, collapse of effective propagation depth, or contraction of representations into low-dimensional subspaces.
1.2  Limitations of Accuracy-Centric Evaluation in Federated GNNs
Machine learning systems are commonly evaluated using task-level metrics such as accuracy, F1-score, or loss. These metrics implicitly assume that improvements or degradation in model performance directly reflect the quality of the learned representations. While this assumption holds reasonably well for centralized learning on i.i.d. data, it becomes insufficient in federated and graph-based learning settings, where model behavior depends critically on relational structure and iterative aggregation dynamics.Primary insight is into accuracy as a Performance Proxy in Conventional Learning Models. In conventional supervised learning, accuracy measures the alignment between model predictions and ground-truth labels on a fixed data distribution. Optimization procedures are designed to minimize empirical risk, and convergence in loss is typically interpreted as convergence in representation quality. Under these assumptions, parameter averaging or gradient-based updates preserve semantic meaning, and degradation in performance is readily observable through scalar metrics.There are some notable limitations of Accuracy Metrics in GNNs.GNNs differ fundamentally from standard models in that predictions are produced through iterative message passing over graph structure. The expressive power of a GNN depends not only on parameter values, but on how those parameters govern information propagation across neighborhoods, graph depth, and connectivity patterns. As a result, a GNN can exhibit stable accuracy while its internal message-passing dynamics degrade.In particular, accuracy does not capture whether:neighborhood information is meaningfully integrated or collapsed,long-range dependencies are preserved or suppressed,node representations remain sensitive to structural perturbations,propagation depth remains effective across layers.These properties are geometric and operator-level in nature, and their degradation may remain invisible to scalar evaluation metrics.Another inference is on accuracy Degradation in Federated Settings.In FL, accuracy is further confounded by the absence of a centralized validation distribution. Each client evaluates performance on its local data, and the global model is rarely assessed on a distribution that reflects the union of client graph structures. As a result, a global model may appear to perform adequately on average while being misaligned with all individual client distributions. Moreover, federated training introduces an iterative feedback loop: the global model is repeatedly broadcast to clients and used as the initialization for local optimization. Even small distortions in model behavior introduced by aggregation can compound over rounds, leading to progressive degradation that is not immediately reflected in early-round accuracy measurements.Empirically, federated GNNs under heterogeneous graph distributions exhibit several failure modes that are poorly explained by accuracy alone:Over-smoothing acceleration: node representations collapse into low-rank subspaces more rapidly than in centralized training.Over-squashing amplification: long-range information is increasingly compressed, reducing the effective receptive field.Update oscillation: local updates increasingly counteract global updates across rounds.Representation drift: the semantic meaning of intermediate embeddings changes across rounds despite stable task loss.These effects indicate degradation of the learned message-passing operator rather than simple underfitting or overfitting.Root Cause can be traced as Aggregation-Induced Geometric Distortion i.e. the common factor underlying these failure modes is the interaction of heterogeneous client updates at the global server.Client updates encode graph-dependent operator perturbations. Standard aggregation mechanisms combine these updates in parameter space without enforcing compatibility of their induced transformations.This process can converge numerically while simultaneously distorting the geometry of the global message-passing operator. Such distortion alters propagation directionality, collapses admissible subspaces, and destabilizes sensitivity to graph structure—effects that directly impair relational learning but may not immediately impact prediction accuracy. These observations indicate that federated GNN training cannot be reliably assessed or stabilized using accuracy-centric metrics alone. Scalar performance measures fail to diagnose degradation of message-passing geometry and provide no insight into the stability of relational representations across communication rounds.Consequently, effective federated GNN systems require mechanisms that explicitly preserve operator-level geometric coherence during aggregation. Without such mechanisms, improvements in optimization or personalization may address symptoms while leaving the underlying structural failure unresolved.
1.3  Problem Formulation and Geometric Failure in Federated GNN Aggregation
1.3.1 Federated GNN Learning Setup
We consider a FL system consisting of $K$ clients and a central server. Each client $k\in\{1,\dots,K\}$ observes a local graph
$\mathcal{G}_{k}=(\mathcal{V}_{k},\mathcal{E}_{k},\mathbf{X}_{k},\mathbf{Y}_{k}),$ (1)
where $\mathcal{V}_{k}$ and $\mathcal{E}_{k}$ denote the node and edge sets, $\mathbf{X}_{k}\in\mathbb{R}^{|\mathcal{V}_{k}|\times d}$ the node features, and $\mathbf{Y}_{k}$ the node labels. Graphs are assumed to be drawn from client-specific distributions $\mathcal{G}_{k}\sim\mathcal{D}_{k}$, with $\mathcal{D}_{k}\neq\mathcal{D}_{k^{\prime}}$ in general.
All clients share a common GNN architecture parameterized by $\theta\in\mathbb{R}^{p}$. Local training at client $k$ minimizes an empirical risk
$\theta_{k}^{(t)}=\arg\min_{\theta}\;\mathbb{E}_{\mathcal{G}_{k}\sim\mathcal{D}_{k}}\left[\mathcal{L}\left(f_{\theta}(\mathcal{G}_{k}),\mathbf{Y}_{k}\right)\right],$ (2)
where $f_{\theta}$ denotes the GNN predictor and $t$ indexes the communication round.
After local optimization, clients transmit parameter updates $\Delta\theta_{k}^{(t)}=\theta_{k}^{(t)}-\theta^{(t)}$ to the server, which computes a global update via aggregation, typically
$\theta^{(t+1)}=\theta^{(t)}+\sum_{k=1}^{K}w_{k}\Delta\theta_{k}^{(t)},$ (3)
where $w_{k}$ are aggregation weights.
1.3.2 GNNs as Parameterized Message-Passing Operators
A $L$-layer GNN can be written as the composition of message-passing operators
$\mathbf{H}^{(l+1)}=\Phi_{\theta^{(l)}}\left(\mathbf{H}^{(l)},\mathcal{G}\right),\quad l=0,\dots,L-1,$ (4)
where $\mathbf{H}^{(0)}=\mathbf{X}$ and $\Phi_{\theta^{(l)}}$ denotes a neighborhood aggregation operator. For instance, in graph convolutional networks,
$\Phi_{\theta}(\mathbf{H},\mathcal{G})=\sigma\left(\tilde{\mathbf{A}}\mathbf{H}\mathbf{W}_{\theta}\right),$ (5)
with normalized adjacency $\tilde{\mathbf{A}}$ and learnable weights $\mathbf{W}_{\theta}$.
Crucially, $\theta$ does not merely parameterize a function mapping features to labels, but induces a linear operator on node representations whose spectral and geometric properties determine information propagation across the graph.
1.3.3 Geometric Representation of Message Passing
Let $\mathcal{T}_{\theta}^{\mathcal{G}}$ denote the message-passing operator induced by parameters $\theta$ on graph $\mathcal{G}$. Linearizing around $\theta$ yields
$\mathcal{T}_{\theta+\delta}^{\mathcal{G}}\approx\mathcal{T}_{\theta}^{\mathcal{G}}+\mathcal{J}_{\theta}^{\mathcal{G}}\delta,$ (6)
where $\mathcal{J}_{\theta}^{\mathcal{G}}$ is the Jacobian of the operator with respect to parameters.
Each client update $\Delta\theta_{k}$ therefore induces a perturbation
$\Delta\mathcal{T}_{k}\;=\;\mathcal{J}_{\theta}^{\mathcal{G}_{k}}\Delta\theta_{k},$ (7)
which lies in an operator space whose geometry depends on $\mathcal{G}_{k}$.
We emphasize that $\mathcal{J}_{\theta}^{\mathcal{G}_{k}}$ differs across clients due to differences in adjacency structure and feature statistics. Consequently, even if $\Delta\theta_{k}$ are comparable in parameter space, their induced operator perturbations $\Delta\mathcal{T}_{k}$ may be geometrically incompatible.
1.3.4 Geometric Incompatibility Under Aggregation
Standard aggregation computes
$\Delta\theta^{\text{global}}=\sum_{k=1}^{K}w_{k}\Delta\theta_{k},$ (8)
which induces a global operator perturbation
$\Delta\mathcal{T}^{\text{global}}=\sum_{k=1}^{K}w_{k}\mathcal{J}_{\theta}^{\mathcal{G}_{k}}\Delta\theta_{k}.$ (9)
Hence,unless the Jacobians $\{\mathcal{J}_{\theta}^{\mathcal{G}_{k}}\}$ share aligned dominant eigenspaces or consistent sensitivity profiles, the resulting perturbation does not correspond to a coherent update of any admissible message-passing operator. Instead, opposing operator directions may cancel or distort one another, leading to a contraction or degeneration of the effective transformation.Geometrically, this corresponds to projecting updates from distinct tangent spaces onto a shared parameter space without enforcing compatibility of their induced operator geometry.We discuss certain consequences for Prediction and Iterative Training which infers,in a GNN relies on repeated application of $\mathcal{T}_{\theta}^{\mathcal{G}}$. When aggregation produces a $\theta^{(t+1)}$ whose induced operator lacks coherent spectral or directional structure, node representations converge toward low-rank or insensitive subspaces. This leads to diminished dependence of predictions on graph structure.Moreover, the federated training process forms a feedback loop. The geometrically distorted $\theta^{(t+1)}$ is broadcast to clients and serves as initialization for the next local optimization phase. Each client must then apply corrective updates to restore operator geometry suitable for its own graph distribution. These corrective updates are again geometrically misaligned across clients, amplifying interference at the next aggregation step and accelerating global operator degradation.Hence we arrive to our problem statement,i.e.,given a FLsystem with heterogeneous graph distributions $\{\mathcal{D}_{k}\}$, standard aggregation fails to preserve the geometric coherence of the induced message-passing operator. In such a case, how global aggregation can be performed such that the induced message-passing operator remains geometrically coherent—preserving directional alignment, subspace structure, and sensitivity across heterogeneous client graphs.
    ormal Derivation of Geometric Mismatch in Federated GNN Aggregation
Figure 2: Horizontal schematic of the GGRS server-side geometric aggregation mechanism.
2.1  Federated Aggregation as Parameter-Space Averaging
Let $\theta\in\mathbb{R}^{p}$ denote the parameters of a fixed GNN architecture. In standard FL, local optimization at client $k$ produces an update
$\Delta\theta_{k}^{(t)}=\theta_{k}^{(t)}-\theta^{(t)},$ (10)
where $\theta^{(t)}$ is the global model at communication round $t$. The server computes the next global model via linear aggregation:
$\theta^{(t+1)}=\theta^{(t)}+\sum_{k=1}^{K}w_{k}\Delta\theta_{k}^{(t)},$ (11)
with aggregation weights $w_{k}\geq 0$, $\sum_{k}w_{k}=1$.
Equation (11) implicitly assumes that the parameter space $\mathbb{R}^{p}$ constitutes a shared, semantically meaningful vector space in which linear combination preserves model behavior.
2.2  GNN Parameters Induce Graph-Dependent Operators
Consider a GNN layer expressed as a message-passing operator
$\mathbf{H}^{(l+1)}=\Phi_{\theta^{(l)}}(\mathbf{H}^{(l)},\mathcal{G}),$ (12)
where $\mathcal{G}$ denotes the underlying graph. Let $\mathcal{T}_{\theta}^{\mathcal{G}}$ denote the composite operator induced by stacking all layers.
Crucially, $\mathcal{T}_{\theta}^{\mathcal{G}}$ depends on both $\theta$ and $\mathcal{G}$. For two distinct graphs $\mathcal{G}_{a}\neq\mathcal{G}_{b}$, the same parameter vector $\theta$ induces distinct operators
$\mathcal{T}_{\theta}^{\mathcal{G}_{a}}\neq\mathcal{T}_{\theta}^{\mathcal{G}_{b}}.$ (13)
Thus, parameters do not correspond to a single global operator, but to a family of graph-dependent operators.
2.3  Local Updates as Operator Perturbations
Let $\theta^{(t)}$ be the global model at round $t$. For client $k$, a local update $\Delta\theta_{k}$ induces a perturbation of the message-passing operator:
$\mathcal{T}_{\theta^{(t)}+\Delta\theta_{k}}^{\mathcal{G}_{k}}\approx\mathcal{T}_{\theta^{(t)}}^{\mathcal{G}_{k}}+\mathcal{J}_{\theta^{(t)}}^{\mathcal{G}_{k}}\Delta\theta_{k},$ (14)
where $\mathcal{J}_{\theta}^{\mathcal{G}_{k}}$ is the Jacobian of the operator with respect to $\theta$ evaluated on graph $\mathcal{G}_{k}$.
The term $\mathcal{J}_{\theta^{(t)}}^{\mathcal{G}_{k}}\Delta\theta_{k}$ represents the directional effect of the update on message passing. Since $\mathcal{J}_{\theta}^{\mathcal{G}_{k}}$ depends on $\mathcal{G}_{k}$, operator perturbations induced by different clients lie in distinct tangent spaces.
2.4  Mismatch Induced by Linear Aggregation
Under FedAvg, the aggregated parameter update is
$\Delta\theta^{\text{global}}=\sum_{k=1}^{K}w_{k}\Delta\theta_{k}.$ (15)
The induced operator perturbation on graph $\mathcal{G}$ is
$\Delta\mathcal{T}^{\mathcal{G}}=\mathcal{J}_{\theta^{(t)}}^{\mathcal{G}}\sum_{k=1}^{K}w_{k}\Delta\theta_{k}.$ (16)
However, the update $\Delta\theta_{k}$ was optimized with respect to $\mathcal{J}_{\theta^{(t)}}^{\mathcal{G}_{k}}$, not $\mathcal{J}_{\theta^{(t)}}^{\mathcal{G}}$. Unless the Jacobians $\{\mathcal{J}_{\theta^{(t)}}^{\mathcal{G}_{k}}\}$ share aligned dominant eigenspaces or consistent sensitivity profiles, Equation (16) does not correspond to a coherent operator update for any $\mathcal{G}$.Geometrically, FedAvg performs linear combination of vectors $\{\Delta\theta_{k}\}$ that originate from incompatible operator tangent spaces, implicitly projecting them onto a shared parameter space without preserving their induced transformation geometry.
2.5  Demonstration of Destructive Interference
Consider two clients $i$ and $j$ whose dominant operator perturbations satisfy
$\langle\mathcal{J}_{\theta}^{\mathcal{G}_{i}}\Delta\theta_{i},\,\mathcal{J}_{\theta}^{\mathcal{G}_{j}}\Delta\theta_{j}\rangle<0.$ (17)
In this case, aggregation induces partial cancellation in the operator space, even if $\|\Delta\theta_{i}\|$ and $\|\Delta\theta_{j}\|$ are individually large. Such cancellation reduces the effective rank or sensitivity of the global operator, leading to contraction of message-passing dynamics.This phenomenon is not captured by parameter norms or loss values, as it arises from misalignment in the induced operator geometry.In the Next section, we discuss Given heterogeneous graph distributions $\{\mathcal{D}_{k}\}$ and locally optimized updates $\{\Delta\theta_{k}\}$, standard federated aggregation constructs a global parameter update that does not, in general, correspond to a geometrically coherent perturbation of the induced message-passing operator.
2.6  Geometric Regulation of Federated GNN Aggregation
Revisiting the Source of Mismatch:
From Section III, local updates $\Delta\theta_{k}$ induce graph-dependent operator perturbations
$\Delta\mathcal{T}_{k}\;=\;\mathcal{J}_{\theta}^{\mathcal{G}_{k}}\Delta\theta_{k},$ (18)
which generally lie in incompatible tangent spaces due to heterogeneity in $\mathcal{G}_{k}$.
We emphasize that the Jacobian-based operator perturbation view in Eqs. (14)–(18) is intended as a first-order conceptual lens for understanding why parameter-space averaging may distort message-passing semantics under heterogeneous client graphs. In practice, the server does not have access to the true client-specific Jacobians $J_{\theta}^{G_{k}}$, nor does it explicitly construct operators in closed form. Rather, the key insight is that local updates are graph-conditioned: the same parameter displacement $\Delta\theta_{k}$ can induce qualitatively different propagation behavior depending on the underlying adjacency structure, feature statistics, and spectral regime of each client. Thus, geometric incompatibility should be interpreted as a mismatch in the directional and sensitivity characteristics of client updates, rather than as an exact computation in an abstract operator manifold. This motivates the use of lightweight server-side proxy measures (such as layerwise directional coherence) to regulate aggregation and preserve stable global message-passing dynamics.Standard aggregation performs linear combination in parameter space, implicitly assuming
$\mathcal{J}_{\theta}^{\mathcal{G}_{k}}\approx\mathcal{J}_{\theta}^{\mathcal{G}_{k^{\prime}}},$ (19)
an assumption that does not hold in federated GNNs.The core objective is therefore not to align parameters directly, but to ensure that the induced operator perturbations remain geometrically compatible under aggregation.
Operator-Space Reformulation of Aggregation:
Let $\mathcal{O}$ denote the space of admissible message-passing operator perturbations. Each client update induces an element $\Delta\mathcal{T}_{k}\in\mathcal{O}$. The server observes $\{\Delta\theta_{k}\}$ but does not have access to $\mathcal{G}_{k}$. However, the effect of each update can be characterized through its action on the current global operator.
We therefore define a server-side mapping
$\Psi:\Delta\theta_{k}\mapsto\widehat{\Delta\mathcal{T}}_{k},$ (20)
where $\widehat{\Delta\mathcal{T}}_{k}$ is a proxy representation of the induced operator perturbation, constructed without access to client data.
Geometric Admissibility Constraints:
We characterize the geometric coherence of operator perturbations through three structural properties: Directional Consistency:Let $\langle\cdot,\cdot\rangle_{\mathcal{O}}$ denote an inner product in operator space. An update $\widehat{\Delta\mathcal{T}}_{k}$ is directionally consistent with the current global operator if
$\langle\widehat{\Delta\mathcal{T}}_{k},\Delta\mathcal{T}^{\text{ref}}\rangle_{\mathcal{O}}\geq 0,$ (21)
where $\Delta\mathcal{T}^{\text{ref}}$ denotes a reference global direction (e.g., a running estimate of admissible operator change). This constraint prevents semantic inversion of message passing under aggregation. Subspace Compatibility:Let $\mathcal{S}\subset\mathcal{O}$ denote a low-dimensional subspace spanned by the dominant propagation modes of the global operator. Each update is projected as
$\Pi_{\mathcal{S}}(\widehat{\Delta\mathcal{T}}_{k}),$ (22)
and aggregation is restricted to this subspace. This preserves diversity of admissible propagation modes while preventing collapse into degenerate directions. Sensitivity Stability:Let $\|\cdot\|_{\mathcal{O}}$ denote a norm capturing sensitivity of the operator to graph perturbations. We impose boundedness:
$\|\widehat{\Delta\mathcal{T}}_{k}\|_{\mathcal{O}}\leq\epsilon,$ (23)
preventing excessive contraction or amplification of message passing.
Geometrically Regulated Aggregation
Under these constraints, server-side aggregation is reformulated as
$\Delta\theta^{\text{global}}=\sum_{k=1}^{K}w_{k}\,\mathcal{R}(\Delta\theta_{k}),$ (24)
where $\mathcal{R}(\cdot)$ denotes a geometric regulation operator that enforces directional consistency, subspace compatibility, and sensitivity stability through the proxy operator representation $\widehat{\Delta\mathcal{T}}_{k}$. Importantly, $\mathcal{R}$ operates solely on transmitted updates and does not require access to client graphs or data.
Preservation of Operator Coherence
Let the regulated operator perturbation be
$\Delta\mathcal{T}^{\text{reg}}=\sum_{k=1}^{K}w_{k}\,\widehat{\Delta\mathcal{T}}_{k}^{\text{reg}}.$ (25)
By construction,
$\Delta\mathcal{T}^{\text{reg}}\in\mathcal{S},\quad\langle\Delta\mathcal{T}^{\text{reg}},\Delta\mathcal{T}^{\text{ref}}\rangle_{\mathcal{O}}\geq 0,\quad\|\Delta\mathcal{T}^{\text{reg}}\|_{\mathcal{O}}\leq\epsilon.$ (26)
Thus, the updated global model
$\theta^{(t+1)}=\theta^{(t)}+\Delta\theta^{\text{global}}$ (27)
induces a message-passing operator whose geometry remains coherent across rounds. Hence,we draw inferences on the effects imposed on Iterative Federated Dynamics.The regulated global model preserves operator geometry.The broadcast model serves as a geometrically aligned initialization for subsequent local optimization.Therefore,Local updates act as refinements within an admissible operator neighborhood,rather than corrective responses to geometric distortion.This breaks the destructive feedback loop identified in previous comprehension and stabilizes federated GNN training under heterogeneous graph distributions.The proposed regulation neither modifies local training nor does it alter the task objective.It does not assume homogeneity across clients as well.Instead, it corrects a structural mismatch by enforcing geometric coherence at the only location where heterogeneous updates interact—the global server.
    ethod
Further, we describe the concrete instantiation of the proposed Global Geometric Reference Structure (GGRS) and its integration into the federated aggregation pipeline. We emphasize that GGRS is a server-side mechanism that operates exclusively on client model updates and does not require access to client data, graph topology, or labels. In FL, the global server is the only entity that simultaneously observes updates from multiple clients. Local clients optimize their models independently and have no visibility into the distribution or interaction of updates produced by other clients. Consequently, any geometric incompatibility arising from the interaction of heterogeneous updates can only be detected and corrected at the server.Furthermore, the global model produced by aggregation serves both as the predictor used for inference and as the initialization for subsequent local optimization rounds. As shown in Section IV, geometric distortion introduced at this stage propagates forward into prediction behavior and backward into local training dynamics. Therefore, preserving geometric coherence at the global server is both necessary and sufficient to stabilize the federated GNN system. GGRS operates by constructing and maintaining a global geometric reference in the space of induced message-passing operator perturbations. At each communication round, this reference serves as an anchor against which incoming client updates are evaluated and regulated.As shown in 3,the construction consists of three components:A proxy representation of operator perturbations induced by client updates;A dynamically maintained reference geometry capturing admissible global propagation behavior and a regulation operator that aligns client updates with the reference prior to aggregation.In cross-domain federation, GGRS regulates and aggregates only the shared encoder parameters, while task-specific classifier heads remain local.
Proxy Representation of Operator Perturbations: Let $\theta^{(t)}$ denote the global model at round $t$, and let $\Delta\theta_{k}^{(t)}$ be the update received from client $k$. As the server does not have access to client graphs, it cannot compute the true operator perturbation $\mathcal{J}_{\theta}^{\mathcal{G}_{k}}\Delta\theta_{k}^{(t)}$. Instead, GGRS constructs a proxy operator representation based on the effect of the update on the current global model.Specifically, the server evaluates the directional effect of $\Delta\theta_{k}^{(t)}$ on the global model parameters and maps it into a low-dimensional geometric summary:
$\mathbf{z}_{k}^{(t)}=\Psi\left(\Delta\theta_{k}^{(t)},\theta^{(t)}\right),$ (28)
where $\Psi(\cdot)$ is a deterministic mapping that captures dominant update directions and relative scaling across layers. In practice, $\Psi$ may be instantiated using layer-wise normalized update projections or low-rank approximations of parameter perturbations. The resulting vector $\mathbf{z}_{k}^{(t)}$ serves as a surrogate for the induced operator perturbation and enables geometric comparison across clients.
Global Geometric Reference Construction: GGRS maintains a reference vector $\mathbf{r}^{(t)}$ that summarizes the admissible direction of global message-passing evolution. This reference is updated incrementally across communication rounds as:
$\mathbf{r}^{(t)}=\alpha\mathbf{r}^{(t-1)}+(1-\alpha)\sum_{k=1}^{K}w_{k}\mathbf{z}_{k}^{(t)},$ (29)
where $\alpha\in[0,1)$ controls the temporal smoothing of the reference and $w_{k}$ are aggregation weights. The reference $\mathbf{r}^{(t)}$ does not represent an optimal update direction for any individual client. Rather, it encodes a geometrically admissible envelope within which client updates can be combined without inducing operator degeneration.
Geometric Regulation of Client Updates: Given the proxy representation $\mathbf{z}_{k}^{(t)}$ and the reference $\mathbf{r}^{(t)}$, GGRS regulates each client update via three operations: 1) Directional Alignment:Updates whose proxy representations exhibit negative alignment with the reference are reflected or downweighted:
$\mathbf{z}_{k}^{\prime}=\begin{cases}\mathbf{z}_{k}^{(t)},&\langle\mathbf{z}_{k}^{(t)},\mathbf{r}^{(t)}\rangle\geq 0,\\ \beta\mathbf{z}_{k}^{(t)},&\text{otherwise},\end{cases}$ (30)
with $\beta\in[0,1)$.Subspace Projection: To prevent collapse into degenerate propagation modes, regulated updates are projected onto a low-dimensional subspace $\mathcal{S}^{(t)}$ spanned by the dominant directions of recent proxy updates:
$\mathbf{z}_{k}^{\prime\prime}=\Pi_{\mathcal{S}^{(t)}}(\mathbf{z}_{k}^{\prime}).$ (31)
Sensitivity Normalization: Finally, updates are rescaled to enforce bounded sensitivity:
$\widehat{\mathbf{z}}_{k}^{(t)}=\frac{\mathbf{z}_{k}^{\prime\prime}}{\max(1,\|\mathbf{z}_{k}^{\prime\prime}\|/\epsilon)}.$ (32)
The regulated proxy $\widehat{\mathbf{z}}_{k}^{(t)}$ is then mapped back to parameter space by applying the same regulation coefficients to $\Delta\theta_{k}^{(t)}$.
GGRS-Enabled Aggregation: The regulated global update is computed as:
$\Delta\theta^{(t)}=\sum_{k=1}^{K}w_{k}\mathcal{R}\left(\Delta\theta_{k}^{(t)}\right),$ (33)
where $\mathcal{R}(\cdot)$ denotes the regulation induced by the proxy operations described above. The global model is updated as:
$\theta^{(t+1)}=\theta^{(t)}+\Delta\theta^{(t)}.$ (34)
We propose a GGRS Algorithm as stated below, and the 3 visualizes the overall GGRS Construction.
Algorithm 1 Federated GNN Training with Global Geometric Regulation (GGRS)
0: Initial global shared encoder $\theta^{(0)}$, number of rounds $T$, clients $\{1,\dots,K\}$, proxy mapping $\Psi(\cdot)$, regulation operator $\mathcal{R}(\cdot)$
1: Initialize geometric reference $\mathbf{r}^{(0)}\leftarrow\mathbf{0}$
2:for $t=0,1,\dots,T-1$ do
3:  Server broadcasts shared encoder $\theta^{(t)}$ to all clients
4:  for all clients $k$ in parallel do
5:   Initialize local encoder with $\theta^{(t)}$ (heads remain local)
6:   Train local GNN on private graph $\mathcal{G}_{k}$
7:   Compute encoder update $\Delta\theta_{k}^{(t)}=\theta_{k}^{(t)}-\theta^{(t)}$
8:   Send $\Delta\theta_{k}^{(t)}$ to server
9:  end for
10:  for all received updates $\Delta\theta_{k}^{(t)}$ do
11:   Compute geometric proxy $\mathbf{z}_{k}^{(t)}=\Psi(\Delta\theta_{k}^{(t)})$
12:   Regulate update $\widehat{\Delta\theta}_{k}^{(t)}=\mathcal{R}(\Delta\theta_{k}^{(t)},\mathbf{z}_{k}^{(t)},\mathbf{r}^{(t)})$
13:  end for
14:  Update reference $\mathbf{r}^{(t+1)}\leftarrow\mathrm{EMA}(\{\mathbf{z}_{k}^{(t)}\},\mathbf{r}^{(t)})$
15:  Aggregate regulated updates $\Delta\theta^{(t)}=\sum_{k=1}^{K}w_{k}\widehat{\Delta\theta}_{k}^{(t)}$
16:  Update global encoder $\theta^{(t+1)}=\theta^{(t)}+\Delta\theta^{(t)}$
17:end for
18:return $\theta^{(T)}$
Figure 3: Construction and operation of the GGRS at the global server.
3.1  Experimental Setup
Experiments are conducted on real-world graph benchmarks that induce heterogeneous message-passing geometry across federated clients, including Amazon Computers and Amazon Photo co-purchase networks [11], as well as the Coauthor-CS collaboration graph [12]. These datasets exhibit substantial variation in topology, feature dimensionality, and class structure, providing a natural testbed for geometric instability in federated aggregation. Amazon Computers contains 13,75213{,}752 nodes and 491,722491{,}722 edges (d=767d{=}767, C=10C{=}10), Amazon Photo contains 7,6507{,}650 nodes and 238,162238{,}162 edges (d=745d{=}745, C=8C{=}8), while Coauthor-CS contains 18,33318{,}333 nodes and 163,788163{,}788 edges with high-dimensional attributes (d=6,805d{=}6{,}805, C=15C{=}15).We adopt a fog–cloud federated graph learning architecture motivated by IoT deployments, where multiple fog-level clients locally optimize private graph objectives and a central cloud server aggregates model updates without access to raw node features, labels, or adjacency. Each client $k$ holds a local graph $G_{k}=(V_{k},E_{k})$ drawn from a domain-specific distribution $P_{k}$, capturing structural and statistical heterogeneity across edge environments.
Federation Regimes. We evaluate the proposed Global Geometric Regulation Scheme (GGRS) under two complementary regimes. First, as the primary benchmark, we consider an intra-domain federated setting in which Amazon Computers is partitioned into $K$ non-IID clients via Dirichlet label-skew sampling, yielding a controlled reference consistent with standard FedAvg, FedSGD, and FedProx baselines. Second, we provide an additional cross-domain stress-test in which distinct datasets (Amazon and Coauthor) are treated as heterogeneous fog silos to highlight geometric degradation under domain shift. In the cross-domain case, only the shared encoder is aggregated while prediction heads remain client-local.Client heterogeneity arises from both structural coupling and feature-induced gradient scale. Graph density,
$\rho=\frac{2|\mathcal{E}|}{|\mathcal{V}|(|\mathcal{V}|-1)},$ (35)
confirms that Amazon graphs operate in dense redundant neighborhood regimes, whereas Coauthor-CS is comparatively sparse. This is further reflected in mean degree $\bar{d}=2|\mathcal{E}|/|\mathcal{V}|$, where dense Amazon clients induce smoother low-frequency message-passing gradients, while Coauthor-CS yields sharper high-variance perturbations amplified by its extreme feature dimension.Each client optimizes a local node classification objective,
$\mathcal{L}_{k}(\theta)=\sum_{v\in\mathcal{V}_{k}^{train}}\ell\big(f_{\theta}(x_{v},\mathcal{N}(v)),y_{v}\big),$ (36)
and transmits parameter displacements $\Delta_{k}^{t}=\theta_{k}^{t}-\theta_{g}^{t}$ after E=1E{=}1 local epoch per round over T=100T{=}100200200 synchronous communication rounds. Under heterogeneous graph spectra, client updates may exhibit weak or negative cosine coherence,
$\gamma_{ij}^{t}=\frac{\langle\Delta_{i}^{t},\Delta_{j}^{t}\rangle}{\|\Delta_{i}^{t}\|\|\Delta_{j}^{t}\|},$ (37)
causing destructive interference under naïve averaging at the cloud server.To resolve this dataset-induced geometric conflict, GGRS operates entirely at the cloud layer by regulating client update vectors with respect to an exponentially smoothed global reference ($\alpha{=}0.9$). Negatively aligned updates are attenuated ($\beta{=}0.5$), ensuring aggregation preserves operator-level geometric coherence despite structural and spectral heterogeneity.
Evaluation Metrics. We report node classification test accuracy alongside geometry-aware stability measures, including mean directional alignment $\Gamma^{t}$ and aggregated update sensitivity norm $S^{t}=\|\Delta^{t}\|_{2}$. All experiments are implemented in PyTorch Geometric and averaged over three random seeds. A toy illustration of geometric regulation is provided in Appendix I.
    esults and Discussion
This section evaluates the proposed Global Geometric Regulation Scheme (GGRS) under heterogeneous federated graph learning and analyzes the geometric stability of global aggregation beyond predictive convergence. All methods follow the synchronous fog–cloud protocol described in Section LABEL:sec:experiments, where $K$ clients locally optimize node classification objectives on private graph partitions and transmit encoder update vectors $\Delta_{k}^{t}$ to a central cloud server for aggregation. In addition to test accuracy, we report geometry-aware diagnostics—directional alignment and global sensitivity norms—which directly capture aggregation coherence of the induced message-passing operator.
Figure 4: Federated training dynamics under heterogeneous client partitions. From left to right: test accuracy, directional alignment, and sensitivity norm across communication rounds. While baseline aggregation converges numerically, GGRS preserves geometric coherence and stabilizes global update structure.
4.1  Geometric Degradation Under Standard Aggregation
[…]
4.2  Comparative Evaluation: FedAvg, FedSGD, FedProx vs. GGRS
We evaluate GGRS by benchmarking it against widely adopted federated baselines: FedAvg , FedSGD [1], and FedProx [li2020federated]. All methods are trained under identical conditions for T=100T=100200200 rounds with $E=1$ local epoch per round. Clients share a common two-layer GCN encoder, and the server aggregates encoder updates with uniform weights. The only difference lies in the aggregation operator: baseline methods apply direct Euclidean averaging, whereas regulated variants apply alignment-aware attenuation prior to aggregation.Across all baselines, GGRS preserves predictive convergence while substantially improving geometric stability. Accuracy trajectories remain comparable, with final differences typically below $1\%$, confirming that regulation does not impair optimization quality. In contrast, geometric coherence differs markedly: regulated aggregation maintains strictly positive alignment throughout training, whereas baseline averaging exhibits late-stage directional collapse.
Figure 5: Late-stage geometric analysis of federated aggregation. Alignment distributions and client-wise directional behavior illustrate that GGRS maintains consistent geometric structure across heterogeneous clients, whereas baseline averaging collapses directionally.
Table 1: Illustrative comparison of baseline aggregation and GGRS after convergence. Values correspond to averages over the final 10 communication rounds (mean $\pm$ std over three seeds).
Method Accuracy (%) $\uparrow$ Alignment $\uparrow$ Sensitivity $\downarrow$
FedAvg 87.8 $\pm$ 0.6 0.05 $\pm$ 0.02 0.11
FedAvg + GGRS 88.3 $\pm$ 0.5 0.29 $\pm$ 0.02 0.10
FedSGD 86.9 $\pm$ 0.8 0.03 $\pm$ 0.02 0.13
FedSGD + GGRS 87.4 $\pm$ 0.7 0.25 $\pm$ 0.03 0.11
FedProx 88.0 $\pm$ 0.5 0.07 $\pm$ 0.03 0.11
FedProx + GGRS 88.5 $\pm$ 0.4 0.31 $\pm$ 0.02 0.10
4.3  Implications for Geometric Regulation
Preserving geometric coherence during aggregation is structurally necessary in federated GNNs. Unlike standard neural models, a GNN parameter vector $\theta$ induces a message-passing operator $\mathcal{T}_{\theta}$ governing neighborhood propagation. Under heterogeneous client graph distributions $\mathcal{G}_{k}\sim\mathcal{D}_{k}$, local updates correspond to graph-dependent operator perturbations that are often directionally incompatible. Standard FedAvg performs linear averaging in Euclidean parameter space without enforcing compatibility of these induced transformations, leading to alignment collapse, sensitivity cancellation, and loss of coherent operator structure even when task accuracy appears stable.The proposed GGRS directly addresses this failure mode by constraining aggregation geometry at the cloud server through alignment-preserving regulation. Importantly, this mechanism is entirely server-side, introduces negligible overhead beyond cosine similarity computation, and is optimizer-agnostic, improving geometric stability across FedAvg, FedSGD, and FedProx without modifying client-side objectives. These results validate that geometric degradation under federated heterogeneity is measurable and correctable only through explicit global regulation.
    onclusion
In this work,we identified a geometric failure mode in heterogeneous federated GNN training that is invisible to conventional accuracy-based evaluation. Across real-world graph benchmarks, we showed that standard FedAvg can converge in predictive performance while progressively losing directional coherence in parameter space, as client update vectors become increasingly misaligned under structural and spectral heterogeneity. To mitigate this pathology, we proposed the GGRS, a lightweight server-side mechanism that monitors update geometry and regulates aggregation to preserve directional consistency and stable global optimization trajectories. From our observation, GGRS maintains substantially higher alignment, reduced variance, and smoother sensitivity behavior than unregulated FedAvg, without sacrificing node classification accuracy. These results establish geometric coherence as a measurable and correctable dimension of federated GNN health.It demonstrates that numerical convergence alone is insufficient under heterogeneity.Further,we would express some extensions which include applying geometric regulation to other structured architectures (e.g., graph transformers), incorporating operator-level summaries such as subspace drift or curvature-aware diagnostics, and developing theoretical proofs on alignment preservation and stability. Further, exploring GGRS to privacy-constrained secure aggregation settings and analyzing its impact on downstream robustness and generalization remain important open directions.
PPENDICES
Appendix A Numerical Spectral Toy Example of Operator Degeneration
In this appendix, we provide a fully numerical graph-based example showing that federated averaging can induce spectral collapse of the global message-passing operator in federated GNNs, even when each client learns a meaningful local propagation regime.
A.1  Setup: One-Layer Linear GCN
Consider a one-layer linear GCN:
$H=\tilde{A}XW,$ (38)
where $\tilde{A}$ is the normalized adjacency operator, $X$ denotes node features, and $W\in\mathbb{R}$ is a scalar propagation weight. For simplicity, we set $X=I$, yielding:
$H=\tilde{A}W.$ (39)
Thus, the learned message-passing operator on client $k$ is:
$T_{k}=\tilde{A}_{k}W_{k}.$ (40)
A.2  Client 1: Chain Graph (Smoothing Regime)
Client 1 holds a 3-node path graph:
123.1-2-3.
The normalized adjacency operator is:
$\tilde{A}_{1}\approx\begin{bmatrix}0.50&0.41&0\\ 0.41&0.33&0.41\\ 0&0.41&0.50\end{bmatrix}$. (41)
Assume local optimization yields a smoothing propagation weight:
$W_{1}=+1,\quad\Rightarrow\quad T_{1}=\tilde{A}_{1}.$ (42)
The eigenvalues of the induced operator are:
$\lambda(T_{1})\approx\{1.00,\;0.50,\;-0.17\}.$ (43)
This corresponds to a stable low-frequency smoothing operator.
A.3  Client 2: Triangle Graph (Opposing Regime)
Client 2 holds a 3-node complete graph:
1231.1-2-3-1.
Its normalized adjacency operator is:
$\tilde{A}_{2}=\frac{1}{3}\begin{bmatrix}1&1&1\\ 1&1&1\\ 1&1&1\end{bmatrix}$. (44)
The eigenvalues satisfy:
$\lambda(\tilde{A}_{2})=\{1,\;0,\;0\}.$ (45)
Assume Client 2 learns an opposing propagation regime:
$W_{2}=-1,\quad\Rightarrow\quad T_{2}=-\tilde{A}_{2}.$ (46)
Thus:
$\lambda(T_{2})=\{-1,\;0,\;0\}.$ (47)
A.4  FedAvg Aggregation Produces Spectral Collapse
Under standard FedAvg:
$W_{\text{avg}}=\frac{1}{2}(W_{1}+W_{2})=\frac{1}{2}(1+(-1))=0.$ (48)
Therefore, the global operator becomes:
$T_{\text{avg}}=\tilde{A}W_{\text{avg}}=0.$ (49)
Hence all eigenvalues collapse:
$\lambda(T_{\text{avg}})=\{0,0,0\}.$ (50)
This implies complete degeneration of message passing: the global model cannot propagate neighborhood information, despite meaningful local operators.
A.5  Geometric Regulation Restores Propagation
GGRS attenuates negatively aligned updates:
$W_{2}^{\prime}=\beta W_{2},\quad\beta=0.5.$ (51)
Thus:
$W_{\text{GGRS}}=\frac{1}{2}(W_{1}+\beta W_{2})=\frac{1}{4}.$ (52)
The regulated global operator becomes:
$T_{\text{GGRS}}=0.25\tilde{A}$, (53)
with eigenvalues:
$\lambda(T_{\text{GGRS}})\approx\{0.25,\;0.125,\;-0.04\}.$ (54)
Thus propagation remains coherent and non-degenerate.Key Insight :This toy example demonstrates that federated averaging may destroy the spectral structure of message passing under heterogeneous client graphs, motivating geometry-aware aggregation mechanisms such as GGRS.demonstrating recovery of geometric coherence:This numeric example illustrates that standard federated averaging can induce geometric degradation—manifested as magnitude collapse and directional inconsistency—even in low-dimensional settings. In contrast, GGRS preserves update magnitude and alignment by explicitly regulating the geometry of aggregated updates. While simplified, this construction mirrors the qualitative behavior observed empirically in heterogeneous federated GNN training.
ACKNOWLEDGMENTS
The authors would like to thank the anonymous reviewers for their constructive feedback, which helped improve the technical clarity and presentation of this manuscript. Consent has been obtained from all individuals acknowledged.
FUNDING INFORMATION
The authors state that no external funding was involved in the conduct of this research.
AUTHOR CONTRIBUTIONS STATEMENT
This journal adopts the Contributor Roles Taxonomy (CRediT) to recognize individual author contributions. All listed authors meet the authorship criteria by contributing to at least one core research activity and one aspect of manuscript preparation. The corresponding author coordinated the research and manuscript submission process.
Table 2: Author Contributions Based on CRediT Taxonomy
[…]
C: Conceptualization M: Methodology So: Software Va: Validation
Fo: Formal Analysis I: Investigation R: Resources D: Data Curation
O: Writing – Original Draft E: Writing – Review & Editing Vi: Visualization Su: Supervision
P: Project Administration Fu: Funding Acquisition
AI ASSISTANCE STATEMENT
AI tools were used only for minor language refinement. All research content and scientific responsibility remain with the authors.
CONFLICT OF INTEREST STATEMENT
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
INFORMED CONSENT
Not applicable. This study does not involve human participants, personal data, or identifiable information.
ETHICAL APPROVAL
Not applicable. This research does not involve human or animal subjects and does not require approval from an ethics committee.
DATA AVAILABILITY
The datasets used in this study are publicly available. Amazon Computers and Amazon Photo datasets are available via the PyTorch Geometric repository, and the Coauthor-CS dataset is publicly accessible. No new datasets were generated during this study.